Given the following data, determine the Chi-Square obtained statistic: a b c d 7 7 10 6 6 1 3 1
Answer #1
Chi-Square Test of independence
Observed Frequencies
0
0 a b c d Total
e 7 7 10 6 30
f 6 1 3 1 11
Total 13 8 13 7 41
Expected frequency of a cell = sum of row*sum of column / total sum
Expected Frequencies
a b c d Total
e 13*30/41=9.512 8*30/41=5.854 13*30/41=9.512 7*30/41=5.122 30
f 13*11/41=3.488 8*11/41=2.146 13*11/41=3.488 7*11/41=1.878 11
Total 13 8 13 7 41
(fo-fe)^2/fe
e 0.6635 0.2245 0.025 0.151
f 1.8095 0.6123 0.068 0.411

Chi-Square Test Statistic,χ² = Σ(fo-fe)^2/fe =   3.9640